OpenStudy (anonymous):

find the exact value of the equation in the interval [0, 2pi) sin3theta-sin6theta=0

OpenStudy (freckles):

you could use double angle identity \[\sin(6 \theta)=\sin(2 \cdot 3 \theta)=2 \sin(3 \theta) \cos(3 \theta)\]

OpenStudy (freckles):

plug and then factor

OpenStudy (anonymous):

how would i factor that ?

OpenStudy (freckles):

both terms have a common factor of sin(3 theta)